The Dawson Company manufactures small lamps and desk lamps. The following shows the activities per product and the total overhead information:

  Setups Inspections Assembly (dlh)
Small Lamps - 3,000 units 8,000 9,000 16,000
Desk Lamps - 6,000 units 16,000 15,000 12,000
       
Activity Pool Activity Base Budgeted Amount
Setups 24,000 $60,000
Inspections 24,000 $120,000
Assembly (dlh) 28,000 $280,000

Calculate the total factory overhead to be charged to desk lamps.
Answer #1
Concepts and reason

Activity-based costing: Activity-based costing is a method used to assign the cost to the products that is apart from traditional costing. The indirect costs are recognized through it. This is mostly used by the manufacturing industries. The cost of products that involve huge materials is transferred to products of fewer contents. Thus, it raises the low volume cost unit.

Fundamentals

Overhead cost: The cost incurred when any equipment is damaged that leads to disassemble is called as overhead cost. It is the cost made to improve or repair the equipment or machinery. It includes the cost of inspection of defective goods.

Cost driver: Cost of the activities is changed using the cost driver. It is the unit of activity. It is a factor that causes change in the activity.

Cost pools: Factory overheads are allocated to production units on the usage of cost pools. Cost of the activities is allocated. It is used in activity-based costing.

Direct labor hours: Direct labor hours are the hours spent by the labor to produce a product directly without other activities.

Calculate the overhead rate:

в
D
Overhead
Rates
А
Activitu peBudgeted
cost
2 Setups $ 60,000
3 Inspections $ 120,000
4 Assembly (dlh) $ 280,000
1
Expected

Therefore, the overhead rate of setups is $2.50, an inspection is $5.00, and for assembly is $10.00.

Working notes:

Calculation is given below:

D
А
Expected use of
Activity Pool Budgeted cost Cost Driver per Overhead Rates
Activity
2 Setups 60000
24000
=B2/C2
3 Inspect

Calculate the total factory overhead to be charged to the desk lamps:

B
C
D
Particulars
Desk lamp Overhead
Desk Lamp
units rate
2 Setups
16,000 $2.50 S 40,000
3 Inspections
15,000 $5.00 $75,000

Therefore, the total factory overhead to be charged to the desk lamps is $235,000.

Working notes:

Calculation is given below:

D
Particulars
Desk Lamp
Desk lamp units Overhead rate
2 Setups 16000
2.5
=B2*C2
3 Inspections 15000
=B3*C3
4 Assembly (dlh) 1

Ans:

The total factory overhead to be charged to the desk lamps is $235,000.